# PlainQAFact: Retrieval-augmented Factual Consistency Evaluation Metric for Biomedical Plain Language Summarization

### v3

@@ Abstract
Hallucinated outputs from large language models (LLMs) pose risks in the medical domain, especially for lay audiences making health-related decisions. Existing automatic factual consistency evaluation methods, such as entailment- and question-answering (QA) -based, struggle with plain language summarization (PLS) due to elaborative explanation phenomenon, which introduces external content (e.g., definitions, background, examples) absent from the scientific abstract to enhance comprehension. To address this, we introduce PlainQAFact, an automatic factual consistency evaluation metric trained on a fine-grained, human-annotated dataset PlainFact, for evaluating factual consistency of both source-simplified and elaborately explained sentences. PlainQAFact first classifies sentence type, then applies a retrieval-augmented QA scoring method. Empirical results show that existing evaluation metrics fail to evaluate the factual consistency in PLS, especially for elaborative explanations, whereas PlainQAFact consistently outperforms them across all evaluation settings. We further analyze PlainQAFact’s effectiveness across external knowledge sources, answer extraction strategies, answer overlap measures, and document granularity levels, refining its overall factual consistency assessment. Taken together, our work presents a sentence-aware, retrieval-augmented metric targeted at elaborative explanations in biomedical PLS tasks, providing the community with both a robust benchmark and a practical tool to advance reliable and safe plain language communication in the medical domain. PlainQAFact and PlainFact are available at: [Uncaptioned image] https://github.com/zhiwenyou103/PlainQAFact.
@@ 1 Introduction
Factual consistency in PLS is typically assessed through a combination of human evaluation and automated metrics [3, 10]. While human evaluation is reliable [11], it is costly and difficult to scale, particularly in biomedical contexts where domain expertise is required. Commonly used factuality metrics can effectively verify content supported by the source but fail to assess factual consistency of added information [12]. However, these metrics depend heavily on high-quality reference summaries, which are often unavailable in plain language summaries. Recent prompt-based evaluation techniques show promise [13, 14], but their sensitivity to factual perturbations in elaborative content remains limited [12].
The lack of suitable benchmark datasets further hinders progress. Many existing datasets are constructed from LLM-generated summaries or apply rule-based perturbations to simulate non-factual content. For example, FactPICO provides expert annotations for plain language summaries of randomized controlled trial abstracts, focusing on PICO elements and evidence inference [9]. However, it includes factuality labels only for added content, leaving simplified sentences unannotated, which are generated by LLMs and potentially inaccurate. In contrast, APPLS perturbs human-written summaries using rule-based transformations [12], but cannot ensure that the resulting outputs remain coherent or factually plausible. These limitations underscore the need for high-quality, sentence-level annotations grounded in human-authored summaries.
@@
To address these challenges, we propose a targeted retrieval-based metric for factual consistency evaluation in PLS (Figure 1). We introduce a new expert-annotated dataset, PlainFact, consisting of human-written plain language summaries aligned with scientific abstracts. Each sentence is labeled with its type (elaborative explanation vs. source simplification), functional role, and alignment to the source (§3.1). Building on this and prior studies for the QA-based evaluation metrics in the general domain [20, 21], we present PlainQAFact, a dual-stage QA-based evaluation metric that selectively applies retrieval only to elaborative explanations (§3.2). Experiment results on several PLS datasets demonstrate the effectiveness of our approach, particularly in assessing complex, elaborative content (§4, §5). Taken together, our work presents the first metric specifically designed to evaluate the factual consistency in elaborative explanations in PLS, providing the community with both a robust benchmark and a practical tool to advance reliable and safe plain language communication in the medical domain.
@@ 2 Related Work
The primary approach for evaluating plain language generation combines automated metrics with human evaluation [3, 10]. While human assessment provides a thorough analysis [11], its high cost and time demands make it impractical for large-scale datasets. Evaluating factual consistency in biomedical PLS is particularly challenging, as it requires domain expertise. Entailment- [15, 16], similarity- [17, 18], model- [19] and QA-based [20, 21] metrics are commonly used for factual consistency assessment but rely heavily on high-quality reference summaries, which are often unavailable or difficult to obtain for PLS. Recent advancements in prompt-based evaluation show promise [13]; however, their sensitivity to factuality perturbations in PLS remains limited [12]. To address these limitations, we propose a reference-free solution for factual consistency evaluation of PLS that effectively assesses factual consistency with external information retrieval to augment the reference summary.
@@
Retrieval-augmented methods enhance text generation by extracting relevant information from external sources to supplement input queries [22]. These methods have been shown to be effective in open-domain QA [23, 24], knowledge-based QA [25], and multi-step reasoning [26, 27]. In the context of PLS, retrieval from structured knowledge bases (KBs) has been shown to improve factual accuracy compared to language models alone [28]. However, retrieval-augmented approaches have not been extensively explored for factual consistency evaluation in PLS, despite their potential for addressing elaborative explanations. In this work, we investigate retrieval-augmented QA to enhance PLS factual consistency assessment while also examining its limitations.
@@ 3 Methods
We first introduce the protocol of PlainFact dataset curation (§3.1). Based on this dataset, we propose PlainQAFact, a two-stage retrieval-augmented QA framework for factual consistency evaluation in PLS tasks (§3.2).
@@ 3.2 PlainQAFact Framework
PlainQAFact conducts fine-grained factual consistency evaluation for plain language summaries by first segmenting each summary into sentences, classifying as either simplification or explanation, and retrieving tailored external knowledge within a retrieval-augmented QA framework. Figure 1 provides an overview of its three key components: sentence-level classification (§3.2.1), domain knowledge retrieval (§3.2.2), and dual-stage QA-based factual consistency evaluation (§3.2.3).
@@ 3.2.3 QA Evaluation Components
QA-based metrics have proved to be effective than other factual consistency evaluation metrics [12] in general summarization tasks and align more closely with human annotations [20, 39]. In this study, we adopt a QA-based approach as a backbone of our evaluation metric to verify factual consistency, incorporating answer extraction, question generation, question filtering, question answering, and answer overlap calculation modules.
@@
The first step in QA-based factual consistency evaluation is to extract answer entities (keyphrases) from plain language summaries as gold answer, then verify factuality by comparing them with answers generated by a QA model for the same questions. If the generated answer is correct or relevant, the summary is considered factual. The previous study used PromptRank [40], a keyphrase generation method based on the T5 model. To compare answer extraction strategies, we use PromptRank as the baseline, and evaluate LLM-based extractors, including an open-source Llama 3.1 8B Instruct model (Llama 3.1) [41] and close-source GPT-4o (GPT-4o) [34]. Prompts for Llama 3.1 and GPT-4o are in E.4.
@@
Questions generated by the QG model (§3.2.3) may not always be answerable. For example, the QG model generates “<What can occur over several days?>,” which is vague and hard to predict the correct answer. To prevent these unanswerable questions from impacting the final evaluation performance, we follow QAFactEval [39] and remove the unanswerable questions using a pre-trained Electra-large model [44]. During QF, the filtering model receives only the plain language summary and its corresponding questions. In the subsequent QA stage (§3.2.3), answers are extracted from the source for answer overlap evaluation, while QF only determines whether the questions can be answered by the plain language summary.
@@ 4.3 Existing Factuality Metrics
To the best of our knowledge, no prior work has efficiently evaluated factuality metrics for detecting elaborative explanations in PLS tasks. Moreover, most factuality metrics are designed for general-domain applications, largely due to limited quality and annotation of existing PLS datasets. To address this gap, we incorporate the following metrics in our experiment, informed by prior work [9]: (1) Dependency-Arc Entailment (DAE) [48] is an entailment-based method that evaluates summary factuality by breaking it into smaller entailment tasks at the arc level. The model independently determines whether each arc’s relationship is supported by the input. (2) AlignScore [19] is a model-based factuality metric using RoBERTa [49]. It extracts claims from the summary and calculates the alignment scores with all the context chunks from the input document. The final score is an average of all highest alignment probabilities. (3) SummaC [16] is an NLI-based inconsistency detection method designed for summarization tasks. It segments documents into sentences and aggregates scores between sentence pairs using a trained NLI model. (4) QAFactEval [39] is a QA-based factuality evaluation metric that assesses summary consistency by generating questions from the source document and comparing the model-generated answers with expected answers. (5) QuestEval [50] is a reference‐less evaluation metric employs a T5 model to generate questions from the source document and verifies whether the summary can correctly answer them.
@@ 5.1 Main Results
In summary, PlainQAFact is a more consistent and effective metric for factual consistency compared to existing automatic approaches. It clearly improves over prior QA-based metrics on CELLS and matches or exceeds them on PlainFact. Although GPT-4o performs better in most of the settings, our PlainQAFact is built entirely on the open-source LLM (i.e., Llama 3.1), ensuring transparency, reproducibility, and accessibility. Its two-step evaluation process, detecting explanations and verifying them with retrieval, helps align metric scores more closely with human judgments in PLS evaluation tasks.
@@ 5.2 Explanation-Only Evaluation
Compared to the results in Table 2, AlignScore is competitive on general PLS tasks but drops on evaluating explanation-only sentences. In contrast, PlainQAFact remains consistently strong when we isolate explanation-only content. PlainQAFact detects added elaborative explanations and checks them with domain knowledge retrieval, which helps when factual consistency depends on facts beyond simple restatement from the source. Since 44% of sentences in our curated benchmark are labeled as elaborative by human annotators, we believe that PlainQAFact is the more suitable and robust factual consistency evaluation metric in this scenario.
@@ 5.3 Ablation Study
Table 5: Error analysis of PlainQAFact with intermediate metric outputs. “Retrieved Knowledge” refers to the source texts retrieved for each plain language sentence. We present correct and failure cases sampled only from elaborative explanation examples (i.e., sentences classified as explanations during evaluation). The correct case has a PlainQAFact score above 0.6, while the failure cases have scores below 0.5.
@@ 6 Discussion and Conclusion
Our study advances the assessment of hallucinations in the field of PLS by supplying both the first domain expert-annotated biomedical PLS benchmark PlainFact and a novel retrieval‐augmented QA‐based factual consistency evaluation metric PlainQAFact. Unlike existing biomedical domain corpora that either lack fine‐grained labels for elaborative content [57] or focus solely on text simplification (e.g., FactPICO[9]), PlainFact captures sentence‐level distinctions, including simplification versus explanation, functional roles, and explicit alignment to source sentences. This granularity not only enables precise error analysis of hallucinations introduced for clarity, but also provides a reusable framework for researchers in other disciplines (e.g., legal, technical) to replicate our annotation protocol and build high‐quality factual consistency datasets tailored to their specialized texts. By offering these data and detailed annotation protocol (A), we anticipate community‐driven extensions, such as multilingual adaptations or integration with domain‐specific ontologies that will democratize development of trustworthy summarization models.
@@
Building on this resource, PlainQAFact differentiates from prior QA‐based metrics (e.g., QAFactEval[21], QuestEval[50]) by first classifying whether a sentence requires external verification given elaborative explanations are often ignored by metrics relying solely on source text, and improve the evaluation accuracy without retrieving for every plain language summary instance. Then, retrieving targeted domain knowledge for each elaborative explanation before posing and answering fact‐checking questions. The ablations demonstrate that selective retrieval yields substantial gains in Kendall’s $\tau$, Pearson correlation, and AUC‐ROC over both end‑to‑end LLM judges[41] and alignment‐based scorers[19], highlighting that judicious incorporation of external evidence is essential for robust factual consistency assessment. Beyond its empirical strengths, PlainQAFact serves as a plug‑and‑play evaluation metric for future summarization systems. Developers can leverage our classifier to classify sentences, apply retrieval only where necessary, thus containing computational costs and obtain interpretable question–answer pairs that manifest non-factual hallucinations. More importantly, PlainQAFact is developed with the open-source backbone model (i.e., Llama 3.1), ensuring broad use with transparency, reproducibility, and lower computational costs. Looking forward, extending our proposed evaluation method to interactive human–AI workflows, where clinicians or subject‐matter experts verify QA outputs on‐the‑fly, could further enhance trust and adoption in high‐stakes domains, including medicine, healthcare, or law.
@@
Despite these promising results, we consider the following limitations of this study. (1) Our error analysis reveals that misclassifications, particularly between source simplifications and elaborative explanations, can lead to retrieval mismatches and unanswerable QA prompts. These issues underscore the need for more nuanced classification and question generation modules. Although we provide substitute solutions for classifying input summaries or sentences, our fine-tuned classifier is limited to the biomedical domain and specifically designed for PLS tasks, fine-tuned through PlainFact. We acknowledge that the dataset used for model fine-tuning is insufficient, resulting in only moderate classification accuracy. Given the scarcity of high-quality human-annotated data for classifying source simplification and elaborative explanation information, future efforts are needed on developing domain-specific (e.g., law, healthcare, social science, etc.) classifiers [56] for factual consistency evaluation. (2) While our experiments indicate that our retrieval-augmented evaluation metric can improve factual consistency assessment in most of the settings, especially in elaborative explanation evaluation, the computational time increases compared to NLI-based evaluation metrics and LLM-based evaluators. The trade-off between evaluation precision and efficiency suggests that further optimization will be beneficial. We suggest to balance wisely based on the trade-offs we report in this study regarding the evaluation time and accuracy. PlainQAFact is a research-oriented metric that prioritizes interpretability and transparency over speed. (3) Given the fair IRA scores, especially for the factuality type field, we acknowledge that our classification method is not fully reliable. The annotated dataset reflects subjective judgments, particularly when distinguishing source simplification from elaborative explanation, and the training set is limited in size. We advise users to apply our pretrained classifier with caution when using it for later factual consistency evaluation steps.

### v4

@@ Abstract
Hallucinated outputs from large language models (LLMs) pose risks in the medical domain, especially for lay audiences making health-related decisions. Existing automatic factual consistency evaluation methods, such as entailment- and question-answering (QA) -based, struggle with plain language summarization (PLS) due to elaborative explanation phenomenon, which introduces external content (e.g., definitions, background, examples) absent from the scientific abstract to enhance comprehension. To address this, we introduce PlainQAFact, an automatic factual consistency evaluation metric trained on a fine-grained, human-annotated dataset PlainFact, for evaluating factual consistency of both source-simplified and elaborately explained sentences. PlainQAFact first classifies sentence type, then applies a retrieval-augmented QA scoring method. Empirical results show that existing evaluation metrics fail to evaluate the factual consistency in PLS, especially for elaborative explanations, whereas PlainQAFact consistently outperforms them across all evaluation settings. We further analyze PlainQAFact’s effectiveness across external knowledge sources, answer extraction strategies, answer overlap measures, and document granularity levels, refining its overall factual consistency assessment. Taken together, our work presents a sentence-aware, retrieval-augmented metric targeted at elaborative explanations in biomedical PLS tasks, providing the community with both a new benchmark and a practical evaluation tool to advance reliable and safe plain language communication in the medical domain. PlainQAFact and PlainFact are available at: [Uncaptioned image] https://github.com/zhiwenyou103/PlainQAFact.
@@ 1 Introduction
Factual consistency in PLS is typically assessed through a combination of human evaluation and automated metrics [3, 10]. While human evaluation is reliable [11], it is costly and difficult to scale, particularly in biomedical contexts where domain expertise is required. Commonly used factual consistency evaluation metrics can effectively verify content supported by the source but fail to assess factual consistency of added information [12]. However, these metrics depend heavily on high-quality reference summaries, which are often unavailable in plain language summaries. Recent prompt-based evaluation techniques show promise [13, 14], but their sensitivity to factual perturbations in elaborative content remains limited [12].
The lack of suitable benchmark datasets further hinders the progress of factual consistency evaluation. Many existing datasets are constructed from LLM-generated summaries or apply rule-based perturbations to simulate non-factual content. For example, FactPICO provides expert annotations for plain language summaries of randomized controlled trial abstracts, focusing on PICO elements and evidence inference [9]. However, it includes factuality labels only for added content, leaving simplified sentences unannotated, which are generated by LLMs and potentially inaccurate. In contrast, APPLS perturbs human-written summaries using rule-based transformations [12], but cannot ensure that the resulting outputs remain coherent or factually plausible. These limitations underscore the need for high-quality, sentence-level annotations grounded in human-authored summaries.
@@
To address these challenges, we first introduce a new expert-annotated dataset, PlainFact, consisting of human-written plain language summaries aligned with scientific abstracts. Each sentence is labeled with its type (elaborative explanation vs. source simplification), functional role, and alignment to the source (§3.1). Building on this and prior studies for the QA-based evaluation metrics in the general domain [20, 21], we propose a targeted retrieval-based metric, PlainQAFact, for factual consistency evaluation in PLS (Figure 1). It is a dual-stage QA-based evaluation metric that selectively applies retrieval only to elaborative explanations (§3.2). Experiment results on several PLS datasets demonstrate the effectiveness of our metric, particularly in assessing complex, elaborative content (§4, §5). Taken together, our work introduces an automatic evaluation metric specifically designed to assess factual consistency for PLS. In addition, we provide the community with a new benchmark that supports nuanced evaluation of simplified and elaborative content. These contributions aim to promote more reliable and safe plain language communication in the medical domain.
@@ 2 Related Work
The primary approach for evaluating plain language generation combines automated metrics with human evaluation [3, 10]. While human assessment provides a thorough analysis [11], its high cost and time demands make it impractical for large-scale datasets. Evaluating factual consistency in biomedical PLS is particularly challenging, as it requires domain expertise. Entailment- [15, 16], similarity- [17, 18], model- [19], and QA-based [20, 21] metrics are commonly used for factual consistency assessment but rely heavily on high-quality reference summaries, which are often unavailable or difficult to obtain for PLS. Recent advancements in prompt-based evaluation show promise [13]; however, their sensitivity to factuality perturbations in PLS remains limited [12]. To address these limitations, we propose a reference-free solution for evaluating factual consistency in PLS, which effectively assesses factual consistency by leveraging external information retrieval to augment the reference summary.
@@
Retrieval-augmented methods enhance text generation by extracting relevant information from external sources to supplement input queries [22]. These methods have been shown to be effective in open-domain QA [23, 24], knowledge-based QA [25], and multi-step reasoning [26, 27]. In the context of PLS, retrieval from structured knowledge bases (KBs) has been shown to improve factual accuracy compared to language models alone [28]. However, retrieval-augmented approaches have not been extensively explored for factual consistency evaluation in PLS, despite their potential for addressing elaborative explanations. In this work, we investigate retrieval-augmented QA scoring to enhance factual consistency assessment in PLS while also examining its limitations.
@@ 3 Methods
We first introduce the protocol of PlainFact dataset curation (§3.1). Based on this dataset, we propose PlainQAFact, a two-stage retrieval-augmented QA scoring framework for factual consistency evaluation in PLS tasks (§3.2).
@@ 3.2 PlainQAFact Framework
PlainQAFact conducts fine-grained factual consistency evaluation for plain language summaries by first segmenting each summary into sentences, classifying as either simplification or explanation, and retrieving tailored external knowledge within a retrieval-augmented QA scoring framework. Figure 1 provides an overview of its three key components: sentence-level classification (§3.2.1), domain knowledge retrieval (§3.2.2), and dual-stage QA-based factual consistency evaluation (§3.2.3).
@@ 3.2.3 QA Evaluation Components
QA-based scoring metrics have proved to be effective than other factual consistency evaluation metrics [12] in general summarization tasks and align more closely with human annotations [20, 39]. In this study, we adopt a QA-based scoring approach as a backbone of our evaluation metric to verify factual consistency, incorporating answer extraction, question generation, question filtering, question answering, and answer overlap calculation modules.
@@
The first step in QA-based factual consistency evaluation is to extract answer entities (keyphrases) from plain language summaries as gold answer, then verify factuality by comparing them with answers generated by a QA model for the same questions. If the generated answer is correct or relevant, the summary is considered factual. The previous study used PromptRank [40], a keyphrase generation method based on the T5 model. To compare answer extraction strategies, we use PromptRank as the baseline, and evaluate LLM-based extractors, including an open-source Llama 3.1 8B Instruct model (Llama 3.1) [41] and close-source GPT-4o (GPT-4o) [34]. Prompts for Llama 3.1 and GPT-4o are presented in E.4.
@@
Questions generated by the QG model (§3.2.3) may not always be answerable. For example, the QG model generates “What can occur over several days?,” which is vague and hard to predict the correct answer. To prevent these unanswerable questions from impacting the final evaluation performance, we follow QAFactEval [39] and remove the unanswerable questions using a pre-trained Electra-large model [44]. During QF, the filtering model receives only the plain language summary and its corresponding questions. In the subsequent QA stage (§3.2.3), answers are extracted from the source for answer overlap evaluation, while QF only determines whether the questions can be answered by the plain language summary.
@@ 4.3 Existing Factuality Metrics
To the best of our knowledge, no prior work has efficiently evaluated factuality metrics for detecting elaborative explanations in PLS tasks. Moreover, most factuality metrics are designed for general-domain applications, largely due to limited quality and annotation of existing PLS datasets. To address this gap, we incorporate the following metrics in our experiment, informed by prior work [9]: (1) Dependency-Arc Entailment (DAE) [48] is an entailment-based method that evaluates summary factuality by breaking it into smaller entailment tasks at the arc level. The model independently determines whether each arc’s relationship is supported by the input. (2) AlignScore [19] is a model-based factuality metric using RoBERTa [49]. It extracts claims from the summary and calculates the alignment scores with all the context chunks from the input document. The final score is an average of all highest alignment probabilities. (3) SummaC [16] is an NLI-based inconsistency detection method designed for summarization tasks. It segments documents into sentences and aggregates scores between sentence pairs using a trained NLI model. (4) QAFactEval [39] is a QA-based factuality evaluation metric that assesses summary consistency by generating questions from the source document and comparing the model-generated answers with expected answers. (5) QuestEval [50] is a reference‐less evaluation metric employs a T5 [51] model to generate questions from the source document and verifies whether the summary can correctly answer them.
@@ 5.1 Main Results
In summary, PlainQAFact is a more consistent and effective metric for factual consistency evaluation compared to existing automatic metrics. It clearly improves over prior QA-based metrics on CELLS and matches or exceeds them on PlainFact. Although GPT-4o performs better in most of the settings, our PlainQAFact is built entirely on the open-source LLM (i.e., Llama 3.1), ensuring transparency, reproducibility, and accessibility. Its two-step evaluation process, detecting explanations and verifying them with retrieval, helps align metric scores more closely with human judgments in PLS evaluation tasks.
@@ 5.2 Explanation-Only Evaluation
Compared to the results in Table 2, AlignScore is competitive on general PLS tasks but drops on evaluating explanation-only sentences. In contrast, PlainQAFact remains consistently strong when we isolate explanation-only content. PlainQAFact detects added elaborative explanations and checks them with domain knowledge retrieval, which helps when factual consistency depends on facts beyond simple restatement from the source. Since 44% of sentences in our curated benchmark are labeled as elaborative explanations by human annotators, we believe that PlainQAFact is the more suitable and robust factual consistency evaluation metric in this scenario.
@@ 5.3 Ablation Study
Table 5: Error analysis of PlainQAFact with intermediate metric outputs. Retrieved Knowledge refers to the source texts retrieved for each plain language sentence. We present correct and failure cases sampled only from elaborative explanation examples (i.e., sentences classified as explanations during evaluation). The correct case has a PlainQAFact score above 0.6, while the failure cases have scores below 0.5.
@@ 6 Discussion and Conclusion
Our study advances the assessment of hallucinations in the field of PLS by supplying both a domain expert-annotated biomedical PLS benchmark PlainFact and a novel retrieval‐augmented QA‐based factual consistency evaluation metric PlainQAFact. Unlike existing biomedical domain corpora that either lack fine‐grained labels for elaborative content [58] or focus solely on text simplification (e.g., FactPICO[9]), PlainFact captures sentence‐level distinctions, including simplification versus explanation, functional roles, and explicit alignment to source sentences. This granularity not only enables precise error analysis of hallucinations introduced for clarity, but also provides a reusable framework for researchers in other disciplines (e.g., legal, technical) to replicate our annotation protocol and build high‐quality factual consistency datasets tailored to their specialized texts. By offering these data and detailed annotation protocol (A), we anticipate community‐driven extensions, such as multilingual adaptations or integration with domain‐specific ontologies that will democratize development of trustworthy summarization models.
@@
Building on this resource, PlainQAFact differentiates from prior QA‐based metrics (e.g., QAFactEval[21], QuestEval[50]) by first classifying whether a sentence requires external verification given elaborative explanations are often ignored by metrics relying solely on source text, and improve the evaluation accuracy without retrieving for every plain language summary instance. Then, retrieving targeted domain knowledge for each elaborative explanation before posing and answering fact‐checking questions. The ablations demonstrate that selective retrieval yields substantial gains in Kendall’s $\tau$, Pearson correlation, and AUC‐ROC over both end‑to‑end LLM judges[41] and alignment‐based scorers[19], highlighting that the incorporation of external evidence is essential for robust factual consistency assessment. Beyond its empirical strengths, PlainQAFact serves as a plug‑and‑play evaluation metric for future summarization systems. Developers can leverage our classifier to classify sentences, apply retrieval only where necessary, thus containing computational costs and obtain interpretable question–answer pairs that manifest non-factual hallucinations. More importantly, PlainQAFact is developed with the open-source backbone model (i.e., Llama 3.1), ensuring broad use with transparency, reproducibility, and lower computational costs. Looking forward, our proposed evaluation metric could serve as a diagnostic tool for researchers and developers to identify and analyze specific factual inconsistencies in PLS systems, guiding future improvements in both model development and evaluation methodology.
@@
Despite these promising results, we consider the following limitations of this study: (1) Our error analysis reveals that misclassifications, particularly between source simplifications and elaborative explanations, can lead to retrieval mismatches and unanswerable QA prompts. These issues underscore the need for more nuanced classification and question generation modules. Given the fair IRA scores, especially for the factuality type field, we acknowledge that our classification method is not fully reliable. The annotated dataset reflects subjective judgments, particularly when distinguishing source simplification from elaborative explanation, and the training set is limited in size. We advise users to apply our pretrained classifier with caution when using it for later factual consistency evaluation steps. Future efforts are needed on developing domain-specific (e.g., law, healthcare, social science, etc.) classifiers [57] for factual consistency evaluation. (2) While our experiments indicate that our retrieval-augmented evaluation metric can improve factual consistency assessment in most of the settings, especially in elaborative explanation evaluation, the computational time increases compared to NLI-based evaluation metrics and LLM-based evaluators. The trade-off between evaluation precision and efficiency suggests that further optimization will be beneficial. We suggest to balance wisely based on the trade-offs we report in this study regarding the evaluation time and accuracy. PlainQAFact is a research-oriented metric that prioritizes interpretability and transparency over speed. (3) The quantitative error analysis on FareBio (Table 11) shows a notable limitation of PlainQAFact in handling numerical content. The answer extraction step may overlook numerical values, preventing them from being verified in subsequent question generation and answering stages. Consequently, PlainQAFact may underestimate factual inconsistency in summaries with incorrect numerical information. (4) PlainFact and PlainQAFact are built upon the CELLS corpus, whose plain language summaries are written by the original article authors. While this ensures domain expertise, it does not guarantee complete factual consistency, as expert-authored summaries may contain unintentional inaccuracies or oversimplifications.